OpenStudy (anonymous):

Determine the value(s) of tan (theta) if cos (theta) = -1/(square root)2 someone can show me how to solve please !!!! Determine the value(s) of tan (theta) if cos (theta) = -1/(square root)2 someone can show me how to solve please !!!! @Mathematics

OpenStudy (anonymous):

cos\[\cos \theta = -1/\sqrt{2} <=> \theta = \cos^{-1} -1/\sqrt{2}, \theta = -\pi/4< => \tan- \pi/4 = -1\]

OpenStudy (anonymous):

how did you get θ=−π/4

OpenStudy (anonymous):

Well you can either memorise a table of all values or work of triangles using SohCahToa. So in this case cos^(-1)-1/sqrt2 I used Cah on a triangle with sides 1 hypotenuse sqrt2. Hard to explain I'm afraid.